1)11,15,18,20,20,23,26,38,42,51

* Find the percentile rank for the score 38

2) If a die is rolled 5 times, find the probability of getting at least one 6.

3)Which score has the best relative position?

A) 3.2 where sample mean is 4.6 and s is 1.5

B) 630 where sample mean is 800 and s is 200

C) 43 where sample mean is 50 and s is 5
Answer #1

1)38 is 8th number

Percentile rank (8/10)×100 %

=80%

2)

Probabilityof getting atleast one six = 1- probability of getting no 6.
Probability of getting no 6 = probability of not getting 6 on first toss*probability of not getting 6 on second toss * .........................* probability of not getting 6 on sixth toss

P = (5/6)*(5/6)*(5/6)*(5/6)*(5/6)*(5/6);

P = 0.6651

3)

A) X1=3.2, M=4.6, S=1.5

Z1=X1-M/S

=3.2-4.6/1.5

=-0.93

B)X2=630

M=800

S=200

Z2=X2-M/S

=630-800/200

=-0.85

C) X3=43, M=50 ,S =5

Z3 =X3-M/S

=43-50/5

= -1.4